# Vision-Language Meets the Skeleton: Progressively Distillation with Cross-Modal Knowledge for 3D Action Representation Learning

### v1

@@ Abstract
Supervised and self-supervised learning are two main training paradigms for skeleton-based human action recognition. However, the former one-hot classification requires labor-intensive predefined action categories annotations, while the latter involves skeleton transformations (e.g., cropping) in the pretext tasks that may impair the skeleton structure. To address these challenges, we introduce a novel skeleton-based training framework (C2VL) based on Cross-modal Contrastive learning that uses the progressive distillation to learn task-agnostic human skeleton action representation from the Vision-Language knowledge prompts. Specifically, we establish the vision-language action concept space through vision-language knowledge prompts generated by pre-trained large multimodal models (LMMs), which enrich the fine-grained details that the skeleton action space lacks. Moreover, we propose the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal contrastive process to progressively control and guide the degree of pulling vision-language knowledge prompts and corresponding skeletons closer. These soft instance discrimination and self-knowledge distillation strategies contribute to the learning of better skeleton-based action representations from the noisy skeleton-vision-language pairs. During the inference phase, our method requires only the skeleton data as the input for action recognition and no longer for vision-language prompts. Extensive experiments show that our method achieves state-of-the-art results on NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD datasets. The code will be available in the future.
@@
Action Recognition, Vision-Language, Cross-Modal, Self-Supervised, Contrastive Learning, Large Multimodal Models.
@@ I Introduction
Action recognition has become a hot topic due to its wide range of application fields, such as human-computer interaction, sports analysis, and dangerous behavior warning [1], etc. In recent years, the continuous development of depth sensors like Kinect [2] has facilitated the easy acquisition of the human body’s skeleton (joints coordinate). Compared with RGB-D sequences, the 3D skeleton is naturally robust to light, background noise, and viewpoint changes while also having the advantages of being lightweight and privacy-preserving.
@@
Skeleton-based action recognition methods are primarily categorized into two distinct paradigms: supervised learning and self-supervised learning. In the former, skeleton data serves as the input, with human-designed action category labels represented as one-hot vectors employed for supervision to design complex network architectures [3, 4, 5]. This training process aims to learn task-specific spatial-temporal representations for actions, which has achieved impressive performance. On the other hand, self-supervised learning approaches generate positive-negative pairs by designing pretext tasks of data transformations (rotation, masking, and cropping, etc.) to explore task-agnostic action representation in a label-free manner based on contrastive learning [6, 7, 8]. These methods leverage pretext tasks for unsupervised pre-training and then finetune for downstream tasks. The results suggest that the self-supervised approaches have the potential to surpass supervised learning methods.
@@
However, while the abovementioned methods have proven effective in representing human actions, several challenges remain. (1) Obtaining numerous labeled data is extremely expensive and labor-intensive for supervised learning. Additionally, the reliance on one-hot vector classification restricts the generality of these methods since additional labeled data is needed to address any other tasks. (2) For self-supervised learning, skeleton data transformations (rotation, masking, cropping, etc.) of the pretext task may impair the structural information inherent in skeleton action, leading to information bias. Inspired by the recent success in multi-modal contrastive learning [9, 10, 11], we consider designing a novel skeleton training paradigm guided by vision-language knowledge to obviate numerous label annotations or compromise the integrity of the skeleton structures, as shown in Fig. 1.
@@
Specifically, our approach leverages sample-specific vision-language knowledge prompts as the supervision to guide the skeleton encoder in learning task-agnostic action representation, which is called C2VL. Firstly, we utilize LMMs (Grounding DINO [12] and LLaVA [13]) as the vision-language engines, employing text prompts and visual questions to generate one-to-one vision knowledge prompts related to human action images and one-to-one language knowledge prompts corresponding to action descriptions, respectively. Subsequently, the vision-language encoders extract features from these knowledge prompts to establish the vision-language action concept space, enriching object-related details and fine-grained descriptions absent in the skeleton action space. However, these one-to-one format skeleton-knowledge pairs exhibit many-to-many correspondences among different samples because LMMs generate similar semantic descriptions for different skeleton actions, which causes the noisy skeleton-vision-language joint action space. Therefore, we introduce the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal self-knowledge distillation learning to progressively control and guide the degree of bringing vision-language knowledge prompts and corresponding skeletons closer in the noisy joint space. This soft contrastive learning process effectively guides the skeleton encoder to learn better task-agnostic 3D human action representation than the hard alignment in CLIP methods. During the inference phase, only skeleton data is needed as the input for action recognition, and vision-language knowledge prompts are omitted without incurring additional computational costs.
@@ II Related Work
II-A Skeleton-based Action Recognition
@@
Supervised skeleton-based action recognition has been the dominant paradigm. CNN-based approaches [14, 15] transform skeleton data into image-like structures, and RNN-based methods [16, 17] delve into long-term contextual relationships from joint sequences, both of them ignore the skeleton’s inherent graph topology. GCN-based methods [3, 18, 4] regard human joints and bones as graph nodes and edges, respectively, enabling effective exploration of the spatial-temporal representation of actions while keeping structural properties. However, the above-mentioned supervised approaches have the disadvantage of requiring extensive human-designed category annotations. Much research pays attention to self-supervised approaches, which are adept at learning task-agnostic representation in a label-free manner. Reconstruction and contrastive learning are two main problem-solving concepts. They reconstruct skeleton structures by masked modeling [19, 20, 21, 22] or generate positive-negative pairs [8, 7, 6, 23, 24] through skeleton transformation (e.g., rotation, cropping) respectively to explore action representation. Unfortunately, these pretext tasks impair the topological structure of the skeleton, leading to information bias. [Summary]: In this paper, we first explore the possibility of learning action representation for skeleton-based action recognition from other modalities (vision and language) without label annotation or skeleton damage.
II-B Multimodal Representation Learning
@@
Numerous multimodal approaches have demonstrated the capacity to learn powerful representations from different modalities. Inspired by their success, foundational vision-language models such as CLIP [9] and ALIGN [25] have gained widespread adoption in the domain of RGB-based action recognition [26, 1, 27, 10, 28, 29, 30, 31, 11]. Many approaches rely on human-designed coarse-grained category names for video action classification [10], pose estimation [28] and action generation [29], etc. Much recent research gradually employs LLMs (GPT-3, etc.) to generate detailed action descriptions of category names for RGB-based action classification [30] and generation [31]. GAP [11] is the first work to utilize LLMs (GPT-3) for skeleton-based action recognition. Regrettably, although these approaches enrich category names with meaningful semantics, the relationship between each action description and skeleton sample falls under the one-to-many format, which belongs to the supervised learning paradigm. [Summary]: In contrast, our proposed framework differs in two aspects. Firstly, we leverage LMMs (Grounding DINO [12] and LLaVA [13]) as offline engines to generate sample-level vision-language knowledge prompts in a one-to-one format without requiring labor-intensive annotations. Second and more importantly, we utilize these one-to-one vision-language knowledge prompts as the supervisory signal to learn skeleton representation in a self-supervised learning paradigm.
II-C Contrastive Learning with Noisy Pairs
@@ III-A Vision-Language Action Concept Space
For a given skeleton dataset $D\in\left\{S_{i}\right\}_{i=1}^{\left|\mathrm{D}\right|}$, we first generate the sample-level vision-language knowledge prompts $\left\{(V_{i},L_{i})\right\}_{i=1}^{\left|\mathrm{D}\right|}$ by offline LMMs (Grounding DINO [12] and LLaVA [13]). These sample-level vision-language knowledge prompts have a one-to-one corresponding with skeleton sequence, which establishes the vision-language action concept space as shown in Fig. 3.
@@ III-A1 Vision Knowledge Prompts
To the best of our knowledge, prior skeleton-based action recognition methods have yet to explore the possibility of utilizing videos to guide the learning of skeleton representation. Here, we introduce the video modality to enrich the motion patterns and object-related details of skeleton representation. These raw videos and skeleton sequences have one-to-one correspondence as they are captured simultaneously by Kinect devices. However, raw videos are challenged to capture fine-grained details and subtle movement changes because of interference from irrelevant background environments. To mitigate this issue, we leverage a pre-trained open-set object detector called Grounding DINO [12], with text prompts “person” to crop the human action images as vision knowledge prompts precisely. These vision knowledge prompts only consist of human action regions enhancing the motion patterns in the vision action concept space.
@@ III-A2 Language Knowledge Prompts
Inspired by the success of CLIP in handling image-text pairs, numerous methods [10, 28, 29, 11] employ human-designed action names and LLMs-generated movement descriptions for action recognition. However, this language information is at the category level and has a one-to-many relationship with the skeleton sequences, so more sample-specific detail descriptions are needed. Therefore, we aim to establish sample-level language action descriptions to enrich the action concept space without requiring expensive annotation and labor-intensive effort. Fortunately, LMMs have an advantage in visual question answering, presenting a new opportunity. Here, we utilize LLaVA [13] with suitable question prompts “Is he/she or are they holding anything in the hand? Is he/she or are they standing or sitting? What is he/she or are they trying to do? Answer the questions concisely” to comprehend and generate action descriptions to the vision action images autonomously. These language knowledge prompts provide sample-specific fine-grained detail descriptions in the language action concept space.
@@ III-C2 Inter-modal Cross-consistency Boosting
where $\mathrm{\mathbf{J}}$ is an all-one matrix. The first term of Equation 7-10 aggregates the information from its cross opposite modality for dual-direction alignment, i.e., alignment from vision $\mathrm{\mathbf{v}}_{i}$ to skeleton $\mathrm{\mathbf{s}}_{j}$ is based on the probability that skeleton $\mathrm{\mathbf{s}}_{i}$ matched with vision $\mathrm{\mathbf{v}}_{j}$. Moreover, the second and third terms are designed to avoid positive pairs that are not sound correspondences for boosting negative pairs with high semantic similarity in cross-modalities.
@@ III-D Progressive Self-distillation Training Objectives
To align skeleton space with vision-knowledge spaces, we adopt a progressive scheme [34] to gradually control and balance the roles of two soft targets throughout training. Specifically, we partition a batch of pairs into two subsets with the refreshed hyper-parameter $\alpha\in[0,1]$ at each epoch. The intra-modal self-similarity soft target is utilized to repair the first subset data $\mathrm{B_{intra}}=\left\lfloor\alpha\mathrm{B}\right\rfloor$ for estimation of the many-to-many relationships, while the inter-class modal cross-consistency is employed to guide the second subset data $\mathrm{B_{inter}}=\mathrm{B}-\left\lfloor\alpha\mathrm{B}\right\rfloor$ for boosting negative pairs with high semantic similarity. At the initial training phase, the value of $\alpha$ should be large for training stability. With the increased training iterations, we gradually decrease the value of $\alpha$ by cosine annealing schedule [42] to increase the contribution of alignment extent between skeleton and high semantic similarity vision-language knowledge prompts. The start value is 0.9, and the end value is 0.1. In this way, the skeleton representation can be better learned in the noisy cross-modal pairs.
@@
In the inference phase, only the skeleton sequences and encoder $E_{s}$ are used for skeleton-based action recognition. The vision-language prompts and respective encoders will be dropped.
@@ IV-A3 PKU-MMD Dataset [45]
TABLE I: Comparison to the state-of-the-art methods for action recognition on the NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD II datasets under the linear evaluation protocol. The best and second-best results are highlighted in bold and underlined, respectively. The J refers to the joint modality, the M denotes the motion modality, and the B is the bone modality.
@@ IV-B Implementation Details
All experiments are conducted using the PyTorch framework on an NVIDIA A100 GPU. Each sample is down-sampled to 64 frames using the data pre-processing code from [11]. We use ST-GCN [52] as our skeleton GCN encoder. For the image-text encoder, we utilize the pre-trained ViT-L/14@336px model from CLIP [9] and freeze its parameters during training. The temperature parameter $\tau$ is 0.07. In the pre-training phase, we train the model with the SGD optimizer for 150 epochs with a batch size 400. The initial learning rate is 0.1 and reduced by a factor of 0.1 at epochs 130 and 140. Weight decay is set to 5e-4 following the strategy in [11]. In the inference stage, we utilize multiple evaluation protocols, including the linear evaluation protocol, finetune protocol, semi-supervised evaluation protocol(1%, 5% or 10% labeled data), KNN evaluation protocol, and transfer learning evaluation protocol for action recognition.

### v2

@@ Abstract
Skeleton-based action representation learning aims to interpret and understand human behaviors by encoding the skeleton sequences, which can be categorized into two primary training paradigms: supervised learning and self-supervised learning. However, the former one-hot classification requires labor-intensive predefined action categories annotations, while the latter involves skeleton transformations (e.g., cropping) in the pretext tasks that may impair the skeleton structure. To address these challenges, we introduce a novel skeleton-based training framework (C2VL) based on Cross-modal Contrastive learning that uses the progressive distillation to learn task-agnostic human skeleton action representation from the Vision-Language knowledge prompts. Specifically, we establish the vision-language action concept space through vision-language knowledge prompts generated by pre-trained large multimodal models (LMMs), which enrich the fine-grained details that the skeleton action space lacks. Moreover, we propose the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal representation learning process to progressively control and guide the degree of pulling vision-language knowledge prompts and corresponding skeletons closer. These soft instance discrimination and self-knowledge distillation strategies contribute to the learning of better skeleton-based action representations from the noisy skeleton-vision-language pairs. During the inference phase, our method requires only the skeleton data as the input for action recognition and no longer for vision-language prompts. Extensive experiments on NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD datasets demonstrate that our method outperforms the previous methods and achieves state-of-the-art results. Code is available at: https://github.com/cseeyangchen/C2VL.
@@
Action Recognition, Vision-Language, Cross-Modal, Self-Supervised, Contrastive Learning, LMMs.
@@ I Introduction
Action recognition has become a hot topic due to its wide range of application fields, such as human-computer interaction, sports analysis, and dangerous behavior warning [1], etc. In the past ten years, the continuous development of depth sensor cameras like Kinect [2] has facilitated the easy acquisition of the human body’s skeleton (3D joints coordinate). Compared with RGB-D sequences, the 3D skeleton is naturally robust to light, background noise, and viewpoint changes while also having the advantages of being lightweight and privacy-preserving.
@@
Skeleton-based action recognition methods are primarily categorized into two distinct paradigms: supervised learning and self-supervised learning. In the former, skeleton data serves as the input, with human-designed action category labels represented as one-hot vectors employed for supervision to design complex network architectures [3, 4, 5]. This training process aims to learn task-specific spatial-temporal representations for actions, which has achieved impressive performance. On the other hand, self-supervised learning approaches generate positive-negative pairs by designing pretext tasks of data transformations (rotation, masking, and cropping, etc.) to explore task-agnostic action representation in a label-free manner based on contrastive learning [6, 7, 8]. These methods leverage pretext tasks for unsupervised pre-training and then fine-tuning them for downstream tasks. The results suggest that the self-supervised approaches have the potential to surpass supervised learning methods.
@@
However, while the aforementioned methods have proven effective in representing human actions, several challenges remain. (1) The one-hot vector classification in supervised learning requires time-consuming and labor-intensive annotating of numerous labels while restricting the generalization performance of these methods. (2) For self-supervised learning, data transformations (rotation, masking, cropping, etc.) of the pretext task may impair the structural information inherent in skeleton action, leading to information bias. For example, the motion of the ”hand rising” action will be impaired if the masking transformation is applied to the hand joints [8]. Inspired by the recent success in multi-modal contrastive learning [9, 10, 11], we consider designing a novel skeleton training paradigm guided by cross-modal representation learning with auto-generated vision-language knowledge prompts, which can boost the task-agnostic generalization capacity while obviating numerous human-effort annotation and irreversible skeleton impairments, as shown in Fig. 1.
@@
Specifically, our approach leverages sample-specific vision-language knowledge prompts as the supervision to guide the skeleton encoder in learning task-agnostic action representation, which is called C2VL. Firstly, we utilize two LMMs (Grounding DINO [12] and LLaVA [13]) as the vision and language engines, respectively. The former is an open-set object detector, and the latter is a foundation language model. Then, we employ text prompts and visual questions for them to auto-generate one-to-one vision knowledge prompts related to human action images and one-to-one language knowledge prompts corresponding to action descriptions, respectively. Subsequently, the vision-language encoders extract features from these knowledge prompts to establish the vision-language action concept space, enriching object-related details and fine-grained descriptions absent in the skeleton action space. However, these one-to-one format skeleton-knowledge pairs exhibit many-to-many correspondences among different samples because LMMs generate similar semantic descriptions for different skeleton actions, which causes the noisy skeleton-vision-language joint action space. Therefore, we introduce the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal self-knowledge distillation learning to progressively control and guide the degree of bringing vision-language knowledge prompts and corresponding skeletons closer in the noisy joint space. This soft contrastive learning process effectively guides the skeleton encoder to learn better task-agnostic 3D human action representation than the hard alignment in CLIP methods. During the inference phase, only skeleton data is needed as the input for action recognition, and vision-language knowledge prompts are omitted without incurring additional computational costs.
@@ II Related Work
II-A Skeleton-based Action Recognition Methods
@@
Supervised skeleton-based action recognition has been the dominant paradigm. CNN-based approaches [14, 15] transform skeleton data into image-like structures, and RNN-based methods [16, 17] delve into long-term contextual relationships from joint sequences, both of them ignore the skeleton’s inherent graph topology. GCN-based methods [3, 18, 4] regard human joints and bones as graph nodes and edges, respectively, enabling effective exploration of the spatial-temporal representation of actions while keeping structural properties. However, the above-mentioned supervised approaches have the disadvantage of requiring extensive human-designed category annotations. Much research pays attention to self-supervised approaches, which are adept at learning task-agnostic representation in a label-free manner. Reconstruction and contrastive learning are two main problem-solving concepts. They reconstruct skeleton structures by masked modeling [19, 20, 21, 22] or generate positive-negative pairs [8, 7, 6, 23, 24] through skeleton transformation (e.g., rotation, cropping) respectively to explore action representation. Unfortunately, these pretext tasks impair the topological structure of the skeleton, leading to information bias [8]. [Summary]: In this paper, we first explore the possibility of learning action representation for skeleton-based action recognition from other modalities (vision and language) without label annotation or skeleton damage.
II-B Multimodal Representation Learning Methods
@@
Numerous multimodal approaches have demonstrated the capacity to learn powerful representations from different modalities [25]. Inspired by their success, foundational vision-language models such as CLIP [9] and ALIGN [26] have gained widespread adoption in the domain of RGB-based action recognition [27, 1, 28, 10, 29, 30, 31, 32, 11]. Many approaches rely on human-designed coarse-grained category names for video action classification [10], pose estimation [29] and action generation [30], etc. PalmCohashNet utilizes multimodal co-learning hash technology for identity authentication [25]. Much recent research gradually employs LLMs (GPT-3, etc.) to generate detailed action descriptions of category names for RGB-based action classification [31] and generation [32]. GAP [11] is the first work to utilize LLMs (GPT-3) for skeleton-based action recognition. Regrettably, although these approaches enrich category names with meaningful semantics, the relationship between each action description and skeleton sample falls under the one-to-many format, which belongs to the supervised learning paradigm. [Summary]: In contrast, our proposed framework differs in two aspects. Firstly, we leverage LMMs (Grounding DINO [12] and LLaVA [13]) as offline engines to generate sample-level vision-language knowledge prompts in a one-to-one format without requiring labor-intensive annotations. Second and more importantly, we utilize these one-to-one vision-language knowledge prompts as the supervisory signal to learn skeleton representation in a self-supervised learning paradigm.
II-C Contrastive Learning with Noisy Pairs Methods
@@ III-A Vision-Language Action Concept Space
For a given skeleton dataset $D\in\left\{S_{i}\right\}_{i=1}^{N}$, we first generate the sample-level vision-language knowledge prompts $\left\{(V_{i},L_{i})\right\}_{i=1}^{N}$ by offline LMMs (Grounding DINO [12] and LLaVA [13]), where $N$ denotes the number of samples in dataset. $S_{i}$ is the $i$-th skeleton sample, $V_{i}$ is the $i$-th vision sample, and $L_{i}$ is the $i$-th language sample. These sample-level vision-language knowledge prompts have a one-to-one corresponding with skeleton sequence, which establishes the vision-language action concept space as shown in Fig. 3.
@@ III-A1 Vision Knowledge Prompts
To the best of our knowledge, prior skeleton-based action recognition methods have yet to explore the possibility of utilizing videos to guide the learning of skeleton representation. Here, we introduce the video modality to enrich the motion patterns and object-related details of skeleton representation. These raw videos and skeleton sequences have one-to-one correspondence as they are captured simultaneously by Kinect devices. However, raw videos are challenged to capture fine-grained details and subtle movements because of interference from irrelevant background environments. To mitigate this issue, we leverage a pre-trained open-set object detector called Grounding DINO [12], with text prompts “person” to crop the human action images as vision knowledge prompts precisely. These vision knowledge prompts only consist of human action regions enhancing the motion patterns in the vision action concept space.
@@ III-A2 Language Knowledge Prompts
Inspired by the success of CLIP in handling image-text pairs, numerous methods [10, 29, 30, 11] employ human-designed action names and LLMs-generated movement descriptions for action recognition. However, this language information is at the category level and has a one-to-many relationship with the skeleton sequences, so more sample-specific detail descriptions are needed. Therefore, we aim to establish sample-level language action descriptions to enrich the action concept space without requiring expensive annotation and labor-intensive effort. Fortunately, LMMs have an advantage in visual question answering, presenting a new opportunity. Here, we utilize LLaVA [13] with suitable question prompts “Is he/she or are they holding anything in the hand? Is he/she or are they standing or sitting? What is he/she or are they trying to do? Answer the questions concisely” to comprehend and generate action descriptions to the vision action images autonomously. These language knowledge prompts provide sample-specific fine-grained descriptions in the language action concept space.
@@ III-C2 Inter-modal Cross-consistency Boosting
where $\mathrm{\mathbf{J}}$ is an all-one matrix. The first term of Equation (7)-(10) aggregates the information from its cross opposite modality for dual-direction alignment, i.e., alignment from vision $\mathrm{\mathbf{v}}_{i}$ to skeleton $\mathrm{\mathbf{s}}_{j}$ is based on the probability that skeleton $\mathrm{\mathbf{s}}_{i}$ matched with vision $\mathrm{\mathbf{v}}_{j}$. Moreover, the second and third terms are designed to avoid positive pairs that are not sound correspondences for boosting negative pairs with high semantic similarity in cross-modalities.
@@ III-D Progressive Self-distillation Training Objectives
To align skeleton space with vision-knowledge spaces, we adopt a progressive scheme [35] to gradually control and balance the roles of two soft targets throughout training. Specifically, we partition a batch of pairs into two subsets with the refreshed hyper-parameter $\alpha\in[0,1]$ at each epoch. The intra-modal self-similarity soft target is utilized to repair the first subset data $\mathrm{B_{intra}}=\left\lfloor\alpha\mathrm{B}\right\rfloor$ for estimation of the many-to-many relationships, while the inter-class modal cross-consistency is employed to guide the second subset $\mathrm{B_{inter}}=\mathrm{B}-\left\lfloor\alpha\mathrm{B}\right\rfloor$ for boosting negative pairs with high semantic similarity. At the initial training phase, the value of $\alpha$ should be large for training stability. With the increased training iterations, we gradually decrease the value of $\alpha$ by cosine annealing schedule [43] to increase the contribution of alignment extent between skeleton and high semantic similarity vision-language knowledge prompts. The start value is 0.9, and the end value is 0.1. In this way, the skeleton representation can be better learned in the noisy cross-modal pairs.
@@
In the inference phase, only the skeleton and encoder $E_{s}$ are used for skeleton-based action recognition. The vision-language prompts and respective encoders will be dropped.
@@ IV-A3 PKU-MMD Dataset [46]
TABLE I: Comparison to the state-of-the-art methods for action recognition accuracy on the NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD II datasets under the linear evaluation protocol. The best and second-best results are highlighted in bold and underlined, respectively. The J refers to the joint modality, the M denotes the motion modality, and the B is the bone modality.
@@ IV-B Implementation Details
All experiments are conducted using the PyTorch framework on an NVIDIA A100 GPU. Each sample is down-sampled to 64 frames using the data pre-processing code from [11]. We use ST-GCN [53] as our skeleton GCN encoder. For the image-text encoder, we utilize the pre-trained ViT-L/14@336px model from CLIP [9] and freeze its parameters during training. The temperature parameter $\tau$ is 0.07. In the pre-training phase, we train the model with the SGD optimizer for 150 epochs with a batch size 400. The initial learning rate is 0.1 and reduced by a factor of 0.1 at epochs 130 and 140. Weight decay is set to 5e-4 following the strategy in [11]. In the inference stage, we utilize multiple evaluation protocols, including the linear evaluation protocol, finetune protocol, semi-supervised evaluation protocol(1%, 5% or 10% labeled data), KNN evaluation protocol, and transfer learning evaluation protocol for action recognition. Meanwhile, following the previous studies [40, 47, 48, 49, 8, 50, 23, 51, 52, 24], several tasks (cross-view, cross-subject, and cross-setup) with different official training-testing split strategies are used to comprehensively validate the algorithm, preventing overfitting in the testing set.